OpenStudy (anonymous):

Hey what is 3x4x6x8x91 and i also need working out P.S is anyone online that doesnt always talk about maths

OpenStudy (apoorvk):

LOoolllLLL the last line :P For your question, use a calculator. (that's a multiply sign, not an \(x\) right?)

OpenStudy (apoorvk):

If that's an \(x\), then use funda: \[\large x^a \times x^b \times....\times x^n = x^{a+b+....+n}\] Use this funda^.

OpenStudy (anonymous):

thanks

OpenStudy (apoorvk):

Lolz no worries.